The highest mark obtained by a student is 10 more than thrice the lowest mark. If the highest mark is 85 find the lowest mark.
Answer by ankor@dixie-net.com(22740)

The highest mark obtained by a student is 10 more than thrice the lowest mark. If the highest mark is 85 find the lowest mark.
let x = the lowest mark
write an equation for exactly what it says
3x + 10 = 85
subtract 10 from both sides
3x = 85 - 10
3x = 75
divide both sides by 3
1x = 75/3
x = 25 is the lowest mark
;
:
See if that is true using the statement
" The highest mark obtained by a student is 10 more than thrice the lowest mark."
85 = 3(25) + 10
85 = 75 + 10